Solve the equation by completing the square. (Enter your answers as a comma-separated list.)
10x2 + 10x + 5 = 0

Answer by stanbon(75887)

Solve the equation by completing the square. (Enter your answers as a comma-separated list.)
10x^2 + 10x + 5 = 0
-------
Divide thru by 5 to get:
2x^2 + 2x + 1 = 0
-------
2x^2 + 2x = -1
-------
Complete the square and maintain the equality::
2(x^2+x+?) = -1 + 2*?
--------------------------
2(x^2 + x + (!/2)^2) = -1 + 2(1/2)^2
------
Factor the left side ; Simplify the right side
2(x+(1/2))^2 = -1 + 1/2
2(x+(1/2)^2 = -1/2
(x+(1/2)^2 = -1/4
-----
Take the square root of both sides to get::
x+(1/2) = i/2 or x+(1/2) = -i/2
x = (-1+i)/2 or x = (-1-i)/2
----------------------------------
Cheers,
Stan H.
-------------